Algebra
the perimeter of a square is 220 feeT? what is the length of each side?
2. A trapezoid has bases 11cm and 4 cm, and a height of 9 cm. What is the area?
3. A rectangular piece of cardboard has a length of 16 inches and an area of 192 square inches. What is the length of the cardboard?

Answer by checkley77(12844)

P=4S
220=4S
4S=220
S=220/4
S=55 ANS. FOR THE LENGTH OF THE SIDES OF THE SQUARE.
-----------------------
A=H(B1+B2)/2
A=9(11+4)/2
A=9*15/2
A=135/2
A=67.5 ANS. FOR THE AREA.
--------------------------
YOU ARE GIVEN THE LENGTH=16.
I BELIEVE THE UNKNOW ITEM IS THE WIDTH.
A=LW
192=16W
16W=192
W=192/16
W=12 ANS. FOR THE WIDTH.